# The LLM Data Auditor: A Metric-oriented Survey on Quality and Trustworthiness in Evaluating Synthetic Data

### v1

@@ Abstract
Large Language Models (LLMs) have emerged as powerful tools for generating data across various modalities. By transforming data from a scarce resource into a controllable asset, LLMs mitigate the bottlenecks imposed by the acquisition costs of real-world data for model training, evaluation, and system iteration. However, ensuring the high quality of LLM-generated synthetic data remains a critical challenge. Existing research primarily focuses on generation methodologies, with limited direct attention to the quality of the resulting data. Furthermore, most studies are restricted to single modalities, lacking a unified perspective across different data types. To bridge this gap, we propose the LLM Data Auditor framework. In this framework, we first describe how LLMs are utilized to generate data across six distinct modalities. More importantly, we systematically categorize intrinsic metrics for evaluating synthetic data from two dimensions: quality and trustworthiness. This approach shifts the focus from extrinsic evaluation, which relies on downstream task performance, to the inherent properties of the data itself. Using this evaluation system, we analyze the experimental evaluations of representative generation methods for each modality and identify substantial deficiencies in current evaluation practices. Based on these findings, we offer concrete recommendations for the community to improve the evaluation of data generation. Finally, the framework outlines methodologies for the practical application of synthetic data across different modalities. Our repository has been released: https://anonymous.4open.science/r/Awesome-LLM-Data-Generation-6457.
@@ 1 Introduction
To ensure that LLM-generated data fulfills its true potential, the research focus must shift from generation techniques to evaluation methodologies. In this survey, we adopt a data-based perspective and establish metrics as our primary organizing principle. Unlike existing surveys oriented toward workflows, lifecycles, or generation reviews (Long et al., 2024; Wang et al., 2024a; Goyal and Mahmoud, 2024), we introduce the LLM Data Auditor framework, as shown in Figure 1 and 2. This framework organizes various data types through a unified structure encompassing 5 core components: LLM-based data generation methods, quality metrics, trustworthy metrics, evaluation gaps, and data usage. Specifically, quality metrics focus on validity, fidelity, and utility to measure fundamental usability, while trustworthy metrics assess safety and faithfulness to identify potential risks. Applying this framework, we analyze the representative literature to identify existing evaluation gaps and discuss practical usages. Ultimately, our framework provides the community with a comprehensive guide for generating and evaluating high-quality, multi-modal data using LLMs.
@@
Agent Data (DTs & Embodied) [Uncaptioned image] Environment & Task Data ChatSUMO Li et al. (2024c), AutoScenario Lu et al. (2024), TTSG Ruan et al. (2025), ODD2CARLA Danso and Büker (2025), SDT-LLM Naeem et al. (2025), L3M+P Agarwal et al. (2025b), SELP Wu et al. (2025), T3 Planner Li and Zhao (2025), PARTNR Chang et al. (2024) Validity ExecRate, $\mathrm{SR}_{\mathrm{valid}}$, PC Fidelity FID/FVD, SSIM, PSNR, LPIPS, $\mathrm{SemFid}$, $\mathrm{MatchRate}$ Diversity AD, RD Utility $\mathrm{SR}_{\mathrm{eval}}$, SimSteps, Offloading, PC, CR, $G_{t}$ Safety RVR, RI, MSCR, $\mathrm{Jerk}_{\mathrm{RMS}}$, HardBrakeRate, SafetySat, Rejection/Risk, minTTC, MDC
Control & Decision Data Grid-Agent Zhang et al. (2025c), Twin-2K-500 Toubia et al. (2025), BEHAVIORCHAIN Li et al. (2025a), LLM Trainer George and Farimani (2025), ELLMER Mon-Williams and others (2025), Instruct2Act Huang et al. (2023), ProgPrompt Singh et al. (2022)
@@ 2.1 Generation Methods
For text data, control is often achieved by processing existing corpora into a trainable mixture with explicit and auditable signals rather than by generating new samples. Modern curation pipelines follow a standard sequence that begins with normalization and language identification. This is followed by the removal of near-duplicate content through approximate matching techniques such as Jaccard or MinHash deduplication across different crawl snapshots. Finally, document-level filtering is performed based on descriptors structured by metadata that are attached to each document (Penedo et al., 2024; 2025). LLMs are increasingly employed as scalable annotators to transform nuanced, abstract attributes—such as educational quality—into operational labels or calibrated scores. By distilling these high-level judgments into lightweight classifiers, researchers can convert subjective criteria into systematic signals. This pipeline enables the efficient filtering and reweighting of massive data mixtures (HuggingFaceFW, 2024).
@@ 2.2 Quality Metrics for Text
Validity requires generations to be structurally well-formed under explicit constraints as well as linguistically acceptable.Beyond structural considerations, we evaluate linguistic well-formedness using a CoLA-style grammatical acceptability scorer. Following prior work, we adopt a sentence-level acceptability classifier to assign acceptability scores to generated sentences (Raman and Shah, 2023). For this purpose, we use a RoBERTa-large model fine-tuned on the CoLA benchmark.Building on this practice, we report a Grammatical Acceptability Rate, also referred to as GAR, as the proportion of generated sentences whose acceptability score exceeds a fixed threshold:
@@
When references are unavailable, we employ USR (Mehri and Eskenazi, 2020), which is a reference free dialogue evaluation framework. This framework provides interpretable sub metrics aligned with five dialogue qualities such as understandability, naturalness, context maintenance, interestingness, and knowledge usage. Following the original formulation, these sub scores are aggregated into an overall quality estimate via a regression model trained to reproduce human overall ratings (Mehri and Eskenazi, 2020). Finally, to detect hallucinations in document grounded generation, we use PMI-FAITH (Nandwani et al., 2023):
@@
$\mathrm{Ent}\text{-}n=-\sum_{g\in\mathcal{G}_{n}}p_{\mathcal{T}}(g)\log p_{\mathcal{T}}(g),$ (11)
@@ 2.5 Usages
More recent approaches integrate LLM based rewards or selectors to identify high value examples. For instance, Jin and Wang (2024) treat a generative LLM as a reward model to score synthetic QA pairs and retain only those predicted to be most useful for training. Survey evidence on data selection for instruction tuning suggests diminishing returns from scaling instruction corpora alone. This motivates data-efficient selection based on utility signals such as quality, diversity, difficulty, and complexity. Empirically, highly curated small sets such as LIMA with 1000 instances and large-scale human AI collaboration corpora such as OpenAssistant conversations can be competitive. Furthermore, automated selectors improve efficiency through model-based filtering and strong LLM scoring using models such as GPT 4 to remove low quality or incorrect examples and retain diverse and high utility instances (Albalak et al., 2024; Köpf et al., 2023).
@@ 3.3 Trustworthy Metrics for Symbolic and Reasoning Data
Higher values indicate that the reasoning progresses through steps supported by semantics and logic instead of abrupt or contradictory jumps. Evaluation methods using entailment have been used to assess the correctness and informativeness of reasoning chains (Prasad et al., 2023). These methods have also been applied to assign partial credit based on textual entailment (Yao and Barbosa, 2024).
@@ 4.1 Generation Methods
A distinct category of methodologies involves fine tuning LLMs specifically for table synthesis. These approaches aim to model structural constraints and distributional dependencies more accurately to enhance both validity and fidelity. For example, GReaT fine tunes an autoregressive language model on serialized tabular rows to enable fully conditional sampling across arbitrary feature subsets. This mechanism improves sample realism and reduces the occurrence of implausible cross feature combinations (Borisov et al., 2023). Extending this architecture, Nguyen et al. (2024) introduce permutation based training combined with feature conditional sampling to preserve feature label correlations and further enhance sample realism. REaLTabFormer targets relational tabular synthesis by generating a parent table autoregressively and conditioning child table generation on the sampled keys (Solatorio and Dupriez, 2023). To address memorization risks and capture inter row structural nuances, HARMONIC incorporates $k$-nearest neighbor-based instructional signals that emphasize neighborhood relations across records (Wang et al., 2024d). Other frameworks integrate explicit self correction mechanisms. Table-LLM-Specialist proposes an iterative generator validator self training paradigm where candidate supervision is generated by models and validated via table specific consistency signals such as permutation and execution invariance. This approach facilitates specialist fine tuning without reliance on manual annotations (Xing et al., 2024). Furthermore, adaptive methods such as TableDreamer progressively synthesize instances that expose model weaknesses by targeting observed failure modes to improve data efficiency and downstream utility (Zheng et al., 2025). Collectively, these approaches directly address challenges identified in recent empirical analyses, specifically constraint and functional dependency violations, by incorporating structural awareness during both training and sampling phases (Xu et al., 2025).
@@ 4.2 Quality Metrics for Tabular Data
Global realism and detectability are assessed by applying a classifier two sample test in the same spirit as TabSyn (Lopez-Paz and Oquab, 2018). This approach utilizes the SDMetrics detection score based on logistic regression (Zhang et al., 2024a; DataCebo, 2024). In this context, we define the area under the curve as the mean receiver operating characteristic area under the curve of the discriminator when distinguishing between real and synthetic data over cross validation splits. SDMetrics defines the detection score as follows DataCebo (2024):
@@ 4.3 Trustworthy Metrics for Tabular Data
We report the differential privacy parameters $\varepsilon$ and $\delta$, which represent the privacy budget, together with downstream utility. This approach allows us to contextualize the trade off between privacy and utility under the standard differential privacy framework (Dwork and Roth, 2014).
@@ 4.4 Evaluation Practice Gap
REaLTabFormer (Solatorio and Dupriez, 2023) $\triangle$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
@@ 4.5 Usage
Synthetic tabular data based on LLMs is increasingly used to facilitate safe data sharing and privacy aware anonymization. By producing records that maintain statistical fidelity to the source distribution while differing from original data points, organizations can release or exchange datasets without compromising sensitive or personally identifiable information. This capability is especially important in regulated domains such as healthcare and finance. In these fields, synthetic data serves as a compliant mechanism for collaboration between different institutions (Miletic and Sariyar, 2024; Barr et al., 2025; Long et al., 2025).
In this context, Miletic and Sariyar (2024) benchmark LLMs based on the Transformer architecture against baseline models such as CTGAN (Xu et al., 2019). They show that larger language models yield superior downstream classification utility and maintain competitive performance even at smaller scales. Furthermore, Barr et al. (2025) show that GPT-4o is capable of generating clinical tabular data in a zero shot manner and achieving high fidelity in terms of the statistical properties within each column. This performance is especially high when the model is guided by descriptive statistics. However, the authors did not directly evaluate downstream utility or the risks of duplication and memorization. They highlighted these areas as important directions for further study.
@@
In situations characterized by data scarcity or a lack of rare feature combinations, generation driven by LLMs provides a way to increase training distributions. This approach helps to improve class balance and the ability of models to generalize to new data (Seedat et al., 2024; Tran and Xiong, 2024; Kim et al., 2025; Yang et al., 2024b).
Recent research is moving beyond simple oversampling methods toward more controllable ways of creating tabular data. In some cases, these methods include formal privacy guarantees. For instance, DP-LLMTGen ensures differential privacy by using a two stage fine tuning pipeline. This process first involves learning the data format and then performing fine tuning with differential privacy using a loss function designed for tabular data, and finally samples the private model to create synthetic tables (Tran and Xiong, 2024).
@@ 5.3.1 Graph Data
In this mathematical framework, $\varepsilon$ represents the privacy budget or privacy loss, which quantifies the maximum allowable change in the output probability distribution when one individual’s information is modified. A smaller $\varepsilon$ value indicates a stronger privacy guarantee because it ensures that the outputs are more similar regardless of whether a specific record is present. The parameter $\delta$ represents the probability that the privacy constraint might be violated, and it is typically required to be a very small value to ensure that the guarantee remains robust (Dwork and Roth, 2014).
When this definition is applied to graph structured data, the concept of neighboring datasets is specialized to either node or edge adjacency (Kasiviswanathan et al., 2013). Node level differential privacy defines neighbors as two graphs where one is obtained from the other by removing a single node along with all its incident edges. In contrast, edge level differential privacy considers graphs to be neighbors if they differ by exactly one edge. Under these adjacency notions, the parameters $\varepsilon$ and $\delta$ serve as quantitative metrics for the risk that a specific user or relationship could be inferred from the generated graph (Kasiviswanathan et al., 2013; Dwork and Roth, 2014).
Similarly, edge-level central differential privacy treats neighboring graphs as those differing by exactly one edge, a concept known as edge adjacency (Kasiviswanathan et al., 2013). Under this adjacency notion, the same $(\varepsilon,\delta)$-differential privacy inequality applies (Dwork and Roth, 2014):
@@
In many local differential privacy settings, one focuses on the pure case described above (Duchi et al., 2014). One can also consider an $(\varepsilon,\delta)$ extension by adding a $\delta$ term as in approximate differential privacy (Dwork and Roth, 2014), but we maintain the common pure form for clarity.
A common implementation in the central model uses global sensitivity and Laplace noise (Dwork and Roth, 2014):
@@
where $f$ is a numeric query, such as degree-histogram bins. The Laplace mechanism yields pure $\varepsilon$-differential privacy, which means $\delta$ is equal to zero, under the chosen adjacency notion (Dwork and Roth, 2014). In graph settings, $\Delta f$ can be very large for degree-related queries under node adjacency, since adding or removing a node can affect many incident edges and consequently many degree counts. Practical approaches therefore impose bounded-degree assumptions or apply degree clipping and projection to a bounded-degree graph family. These techniques keep $\Delta f$ and the required noise level manageable (Kasiviswanathan et al., 2013).
In our evaluation, the privacy parameters $\varepsilon$ and $\delta$ for central differential privacy, and $\varepsilon$ for pure local differential privacy, serve as quantitative privacy metrics. A smaller $\varepsilon$ and a smaller $\delta$ indicate stronger privacy protection, which typically comes at the cost of utility (Dwork and Roth, 2014).
Finally, it is common to require $\delta$ to be negligible in an appropriate problem or security parameter. It is important to avoid regimes where $\delta$ is non-trivially large, such as on the order of one over $n$ for population size $n$ (Dwork and Roth, 2014). In graph settings where nodes correspond to protected individuals, a practical heuristic is to select $\delta$ to be far smaller than one over the number of vertices in the graph when using approximate differential privacy. The exact choice remains dependent on the specific application and threat model.

### v2

# The LLM Data Auditor A Metric-oriented Survey on Quality and Trustworthiness in Evaluating Synthetic Data

@@ Abstract
Large Language Models (LLMs) have emerged as powerful tools for generating data across various modalities. By transforming data from a scarce resource into a controllable asset, LLMs mitigate the bottlenecks imposed by the acquisition costs of real-world data for model training, evaluation, and system iteration. However, ensuring the high quality of LLM-generated synthetic data remains a critical challenge. Existing research primarily focuses on generation methodologies, with limited direct attention to the quality of the resulting data. Furthermore, most studies are restricted to single modalities, lacking a unified perspective across different data types. To bridge this gap, we propose the LLM Data Auditor framework. In this framework, we first describe how LLMs are utilized to generate data across six distinct modalities. More importantly, we systematically categorize intrinsic metrics for evaluating synthetic data from two dimensions: quality and trustworthiness. This approach shifts the focus from extrinsic evaluation, which relies on downstream task performance, to the inherent properties of the data itself. Using this evaluation system, we analyze the experimental evaluations of representative generation methods for each modality and identify substantial deficiencies in current evaluation practices. Based on these findings, we offer concrete recommendations for the community to improve the evaluation of data generation. Finally, the framework outlines methodologies for the practical application of synthetic data across different modalities.
[…]
@@ 1 Introduction
To ensure that LLM-generated data fulfills its true potential, the research focus must shift from generation techniques to evaluation methodologies. In this survey, we adopt a data-based perspective and establish metrics as our primary organizing principle. Unlike existing surveys oriented toward workflows, lifecycles, or generation reviews (Long et al., 2024; Wang et al., 2024a; Goyal & Mahmoud, 2024), we introduce the LLM Data Auditor framework, as shown in Figure 1 and 2. This framework organizes various data types through a unified structure encompassing 5 core components: LLM-based data generation methods, quality metrics, trustworthy metrics, evaluation gaps, and data usage. Specifically, quality metrics focus on validity, fidelity, and utility to measure fundamental usability, while trustworthy metrics assess safety and faithfulness to identify potential risks. Applying this framework, we analyze the representative literature to identify existing evaluation gaps and discuss practical usages. Ultimately, our framework provides the community with a comprehensive guide for generating and evaluating high-quality, multi-modal data using LLMs.
@@
Agent Data (DTs & Embodied) [Uncaptioned image] Environment & Task Data ChatSUMO Li et al. (2024c), AutoScenario Lu et al. (2024), TTSG Ruan et al. (2025), ODD2CARLA Danso & Büker (2025), SDT-LLM Naeem et al. (2025), L3M+P Agarwal et al. (2025b), SELP Wu et al. (2025), T3 Planner Li & Zhao (2025), PARTNR Chang et al. (2024) Validity ExecRate, $\mathrm{SR}_{\mathrm{valid}}$, PC Fidelity FID/FVD, SSIM, PSNR, LPIPS, $\mathrm{SemFid}$, $\mathrm{MatchRate}$ Diversity AD, RD Utility $\mathrm{SR}_{\mathrm{eval}}$, SimSteps, Offloading, PC, CR, $G_{t}$ Safety RVR, RI, MSCR, $\mathrm{Jerk}_{\mathrm{RMS}}$, HardBrakeRate, SafetySat, Rejection/Risk, minTTC, MDC
Control & Decision Data Grid-Agent Zhang et al. (2025c), Twin-2K-500 Toubia et al. (2025), BEHAVIORCHAIN Li et al. (2025a), LLM Trainer George & Farimani (2025), ELLMER Mon-Williams et al. (2025), Instruct2Act Huang et al. (2023), ProgPrompt Singh et al. (2022)
@@ 2.1 Generation Methods
For text data, control is often achieved by processing existing corpora into a trainable mixture with explicit and auditable signals rather than by generating new samples. Modern curation pipelines follow a standard sequence that begins with normalization and language identification. This is followed by the removal of near-duplicate content through approximate matching techniques such as Jaccard or MinHash deduplication across different crawl snapshots. Finally, document-level filtering is performed based on descriptors structured by metadata that are attached to each document (Penedo et al., 2024, 2025). LLMs are increasingly employed as scalable annotators to transform nuanced, abstract attributes—such as educational quality—into operational labels or calibrated scores. By distilling these high-level judgments into lightweight classifiers, researchers can convert subjective criteria into systematic signals. This pipeline enables the efficient filtering and reweighting of massive data mixtures (HuggingFaceFW, 2024).
@@ 2.2 Quality Metrics for Text
Validity requires generations to be structurally well-formed under explicit constraints as well as linguistically acceptable.Beyond structural considerations, we evaluate linguistic well-formedness using a CoLA-style grammatical acceptability scorer. Following prior work, we adopt a sentence-level acceptability classifier to assign acceptability scores to generated sentences (Raman & Shah, 2023). For this purpose, we use a RoBERTa-large model fine-tuned on the CoLA benchmark.Building on this practice, we report a Grammatical Acceptability Rate, also referred to as GAR, as the proportion of generated sentences whose acceptability score exceeds a fixed threshold:
@@
When references are unavailable, we employ USR (Mehri & Eskenazi, 2020), which is a reference free dialogue evaluation framework. This framework provides interpretable sub metrics aligned with five dialogue qualities such as understandability, naturalness, context maintenance, interestingness, and knowledge usage. Following the original formulation, these sub scores are aggregated into an overall quality estimate via a regression model trained to reproduce human overall ratings (Mehri & Eskenazi, 2020). Finally, to detect hallucinations in document grounded generation, we use PMI-FAITH (Nandwani et al., 2023):
@@
$\mathrm{Ent}\text{-}n=-\sum_{g\in\mathcal{G}_{n}}p_{\mathcal{T}}(g)\log p_{\mathcal{T}}(g),$
@@ 2.5 Usages
More recent approaches integrate LLM based rewards or selectors to identify high value examples. For instance, Jin & Wang (2024) treat a generative LLM as a reward model to score synthetic QA pairs and retain only those predicted to be most useful for training. Survey evidence on data selection for instruction tuning suggests diminishing returns from scaling instruction corpora alone. This motivates data-efficient selection based on utility signals such as quality, diversity, difficulty, and complexity. Empirically, highly curated small sets such as LIMA with 1000 instances and large-scale human AI collaboration corpora such as OpenAssistant conversations can be competitive. Furthermore, automated selectors improve efficiency through model-based filtering and strong LLM scoring using models such as GPT 4 to remove low quality or incorrect examples and retain diverse and high utility instances (Albalak et al., 2024; Köpf et al., 2023).
@@ 3.3 Trustworthy Metrics for Symbolic and Reasoning Data
Higher values indicate that the reasoning progresses through steps supported by semantics and logic instead of abrupt or contradictory jumps. Evaluation methods using entailment have been used to assess the correctness and informativeness of reasoning chains (Prasad et al., 2023). These methods have also been applied to assign partial credit based on textual entailment (Yao & Barbosa, 2024).
@@ 4.1 Generation Methods
A distinct category of methodologies involves fine tuning LLMs specifically for table synthesis. These approaches aim to model structural constraints and distributional dependencies more accurately to enhance both validity and fidelity. For example, GReaT fine tunes an autoregressive language model on serialized tabular rows to enable fully conditional sampling across arbitrary feature subsets. This mechanism improves sample realism and reduces the occurrence of implausible cross feature combinations (Borisov et al., 2023). Extending this architecture, Nguyen et al. (2024) introduce permutation based training combined with feature conditional sampling to preserve feature label correlations and further enhance sample realism. REaLTabFormer targets relational tabular synthesis by generating a parent table autoregressively and conditioning child table generation on the sampled keys (Solatorio & Dupriez, 2023). To address memorization risks and capture inter row structural nuances, HARMONIC incorporates $k$-nearest neighbor-based instructional signals that emphasize neighborhood relations across records (Wang et al., 2024d). Other frameworks integrate explicit self correction mechanisms. Table-LLM-Specialist proposes an iterative generator validator self training paradigm where candidate supervision is generated by models and validated via table specific consistency signals such as permutation and execution invariance. This approach facilitates specialist fine tuning without reliance on manual annotations (Xing et al., 2024). Furthermore, adaptive methods such as TableDreamer progressively synthesize instances that expose model weaknesses by targeting observed failure modes to improve data efficiency and downstream utility (Zheng et al., 2025). Collectively, these approaches directly address challenges identified in recent empirical analyses, specifically constraint and functional dependency violations, by incorporating structural awareness during both training and sampling phases (Xu et al., 2025).
@@ 4.2 Quality Metrics for Tabular Data
Global realism and detectability are assessed by applying a classifier two sample test in the same spirit as TabSyn (Lopez-Paz & Oquab, 2018). This approach utilizes the SDMetrics detection score based on logistic regression (Zhang et al., 2024a; DataCebo, 2024). In this context, we define the area under the curve as the mean receiver operating characteristic area under the curve of the discriminator when distinguishing between real and synthetic data over cross validation splits. SDMetrics defines the detection score as follows DataCebo (2024):
@@ 4.3 Trustworthy Metrics for Tabular Data
We report the differential privacy parameters $\varepsilon$ and $\delta$, which represent the privacy budget, together with downstream utility. This approach allows us to contextualize the trade off between privacy and utility under the standard differential privacy framework (Dwork & Roth, 2014).
@@ 4.4 Evaluation Practice Gap
REaLTabFormer (Solatorio & Dupriez, 2023) $\triangle$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
@@ 4.5 Usage
Synthetic tabular data based on LLMs is increasingly used to facilitate safe data sharing and privacy aware anonymization. By producing records that maintain statistical fidelity to the source distribution while differing from original data points, organizations can release or exchange datasets without compromising sensitive or personally identifiable information. This capability is especially important in regulated domains such as healthcare and finance. In these fields, synthetic data serves as a compliant mechanism for collaboration between different institutions (Miletic & Sariyar, 2024; Barr et al., 2025; Long et al., 2025).
In this context, Miletic & Sariyar (2024) benchmark LLMs based on the Transformer architecture against baseline models such as CTGAN (Xu et al., 2019). They show that larger language models yield superior downstream classification utility and maintain competitive performance even at smaller scales. Furthermore, Barr et al. (2025) show that GPT-4o is capable of generating clinical tabular data in a zero shot manner and achieving high fidelity in terms of the statistical properties within each column. This performance is especially high when the model is guided by descriptive statistics. However, the authors did not directly evaluate downstream utility or the risks of duplication and memorization. They highlighted these areas as important directions for further study.
@@
In situations characterized by data scarcity or a lack of rare feature combinations, generation driven by LLMs provides a way to increase training distributions. This approach helps to improve class balance and the ability of models to generalize to new data (Seedat et al., 2024; Tran & Xiong, 2024; Kim et al., 2025; Yang et al., 2024b).
Recent research is moving beyond simple oversampling methods toward more controllable ways of creating tabular data. In some cases, these methods include formal privacy guarantees. For instance, DP-LLMTGen ensures differential privacy by using a two stage fine tuning pipeline. This process first involves learning the data format and then performing fine tuning with differential privacy using a loss function designed for tabular data, and finally samples the private model to create synthetic tables (Tran & Xiong, 2024).
@@ 5.3.1 Graph Data
In this mathematical framework, $\varepsilon$ represents the privacy budget or privacy loss, which quantifies the maximum allowable change in the output probability distribution when one individual’s information is modified. A smaller $\varepsilon$ value indicates a stronger privacy guarantee because it ensures that the outputs are more similar regardless of whether a specific record is present. The parameter $\delta$ represents the probability that the privacy constraint might be violated, and it is typically required to be a very small value to ensure that the guarantee remains robust (Dwork & Roth, 2014).
When this definition is applied to graph structured data, the concept of neighboring datasets is specialized to either node or edge adjacency (Kasiviswanathan et al., 2013). Node level differential privacy defines neighbors as two graphs where one is obtained from the other by removing a single node along with all its incident edges. In contrast, edge level differential privacy considers graphs to be neighbors if they differ by exactly one edge. Under these adjacency notions, the parameters $\varepsilon$ and $\delta$ serve as quantitative metrics for the risk that a specific user or relationship could be inferred from the generated graph (Kasiviswanathan et al., 2013; Dwork & Roth, 2014).
Similarly, edge-level central differential privacy treats neighboring graphs as those differing by exactly one edge, a concept known as edge adjacency (Kasiviswanathan et al., 2013). Under this adjacency notion, the same $(\varepsilon,\delta)$-differential privacy inequality applies (Dwork & Roth, 2014):
@@
In many local differential privacy settings, one focuses on the pure case described above (Duchi et al., 2014). One can also consider an $(\varepsilon,\delta)$ extension by adding a $\delta$ term as in approximate differential privacy (Dwork & Roth, 2014), but we maintain the common pure form for clarity.
A common implementation in the central model uses global sensitivity and Laplace noise (Dwork & Roth, 2014):
@@
where $f$ is a numeric query, such as degree-histogram bins. The Laplace mechanism yields pure $\varepsilon$-differential privacy, which means $\delta$ is equal to zero, under the chosen adjacency notion (Dwork & Roth, 2014). In graph settings, $\Delta f$ can be very large for degree-related queries under node adjacency, since adding or removing a node can affect many incident edges and consequently many degree counts. Practical approaches therefore impose bounded-degree assumptions or apply degree clipping and projection to a bounded-degree graph family. These techniques keep $\Delta f$ and the required noise level manageable (Kasiviswanathan et al., 2013).
In our evaluation, the privacy parameters $\varepsilon$ and $\delta$ for central differential privacy, and $\varepsilon$ for pure local differential privacy, serve as quantitative privacy metrics. A smaller $\varepsilon$ and a smaller $\delta$ indicate stronger privacy protection, which typically comes at the cost of utility (Dwork & Roth, 2014).
Finally, it is common to require $\delta$ to be negligible in an appropriate problem or security parameter. It is important to avoid regimes where $\delta$ is non-trivially large, such as on the order of one over $n$ for population size $n$ (Dwork & Roth, 2014). In graph settings where nodes correspond to protected individuals, a practical heuristic is to select $\delta$ to be far smaller than one over the number of vertices in the graph when using approximate differential privacy. The exact choice remains dependent on the specific application and threat model.